# Super heavy dark matter origin of the PeV neutrino event: KM3-230213A

### v1

@@ V Stochastic gravitational waves from disappearing domain walls
BPGW1 $1.47\times 10^{5}$$9.73\times 10^{5}$ 0.221.14×$1.14\times 10^{-2}$$9.28\times 10^{7}$
[…]
VI Conclusions and future directions
We explore the possibility that the recently observed high-energy neutrino event, KM3-230213A, reported by the KM3NeT collaboration, originates from the decay of super-heavy dark matter (SHDM). We consider an extension of the Standard Model that includes a right-handed neutrino ($N$), a singlet scalar ($S$), and a singlet fermion ($\chi$), which serves as the DM candidate. The model is supplemented by a discrete $Z_{2}$ symmetry, under which $S$ and $\chi$ are odd, while all other particles remain even. This $Z_{2}$ symmetry is spontaneously broken when $S$ acquires a vacuum expectation value (vev), leading to a mixing between $\chi$ and $N$. Consequently, the DM can decay into neutrinos. We constrain the DM lifetime using IceCube’s upper limits, along with the KM3NeT observation of KM3-230213A. Given the large uncertainties in the observed flux, we find that the event can be explained while remaining consistent with IceCube’s upper bound for a DM mass of $4.5\times 10^{8}$ GeV and a lifetime of $5\times 10^{29}$ s. We place limits on the lifetime of SHDM as a function of its mass. In our model, DM can also produce gamma rays. However, we find that current gamma-ray constraints are significantly weaker than the predicted flux for the same parameters that explain the KM3NeT observation, ensuring the viability of our framework. Furthermore, the breaking of the $Z_{2}$ symmetry leads to the formation of domain walls, whose subsequent collapse can generate a stochastic gravitational wave background. We estimated the expected GW spectrum and discussed its detectability in future observatories. The presence of such a GW signal would provide an independent probe of the model, offering a multi-messenger approach to studying SHDM through neutrinos, gamma rays, and gravitational waves.
If the reheating temperature ($T_{\rm RH}$) of the Universe is less than the mass of the DM, the DM relic might be produced through the dilute plasma or some non-thermal processes. One of the typical scenarios would be that the DM relic can be directly produced by the inflaton ($\phi$) decay. In this case, the abundance of the DM can be given by, e.g., $Y_{\rm DM}=(3/4){\rm Br}~{}T_{\rm RH}/M_{\phi}$ where $M_{\phi}$ is the mass of the inflation and $\rm Br$ is the branching fraction of $\phi$ decay to the DM.
Acknowledgements.
K.K. is supported by KAKENHI Grant No. JP23KF0289, No. JP24H01825, and No. JP24K07027. P.K.P. would like to acknowledge the Ministry of Education, Government of India, for providing financial support for his research via the Prime Minister’s Research Fellowship (PMRF) scheme. The work of N.S. is supported by the Department of Atomic Energy Board of Research in Nuclear Sciences, Government of India (Ref. Number: 58/14/15/2021-BRNS/37220).

### v2

@@ Appendix A Alleviating the tension with IceCube’s and Auger’s null detection
where we use the effective areas from Refs. [1, 65, 66]. The observation times are set as 335 days, 9 years, and 18 years for KM3NeT, IceCube, and Auger, respectively.
We now construct the likelihood for observing $N^{i}_{\rm obs}$ events given an expectation $N^{i}_{\rm exp}$ from SHDM decay as
$\mathcal{L}_{i}=\frac{(N^{i}_{\rm exp})^{N^{i}_{\rm obs}}e^{-N^{i}_{\rm exp}}}{N^{i}_{\rm obs}!}$ (9)
The joint likelihood over all detectors is
$\mathcal{L}_{\rm total}=\prod_{i}\mathcal{L}_{i}$ (10)
We then compute the log-likelihood as
$\log\mathcal{L}_{\rm total}=\sum_{i}\bigg{(}N^{i}_{\rm obs}\log N^{i}_{\rm exp}-N^{i}_{\rm exp}\bigg{)}.$ (11)
The test statistics (TS) is defined as
${\rm TS}=-2(\log\mathcal{L}_{\rm total}-\log\mathcal{L}_{\rm max}),$ (12)
where $\log\mathcal{L}_{\rm max}$ is the maximum log-likelihood. Under Wilks’ theorem, TS follows a chi-squared distribution with one degree of freedom ( which is the lifetime of the DM in our case), allowing us to translate it into a Gaussian significance (in units of standard deviations) via
$\sigma=\sqrt{\rm TS}$ (13)
Figure 4: [Left]: Significance of the SHDM interpretation of the KM3-230213A event as a function of the dark matter lifetime $\tau_{\rm DM}$, for representative masses $M_{\rm DM}=1.5\times 10^{8},4.5\times 10^{8},5.2\times 10^{9}$ GeV. The significance quantifies the tension between the SHDM predicted neutrino flux and the non-observation. [Right]: Neutrino fluxes from SHDM decay for $M_{\rm DM}=4.5\times 10^{8}$ GeV and $\tau_{\rm DM}=3.46\times 10^{29}$ s ($\tau_{\rm DM}=1.2\times 10^{30}$ s) represented by light blue solid line (blue dashed line) compared with upper limits from IceCube/EHE and Auger, as well as measurements from IceCube-HESE, NST, and Glashow events. The red point marks the KM3-230213A event.
The SHDM decay offers a viable explanation that can reproduce the observed flux while remaining compatible with existing null results. As shown in the left panel of Fig. 4, by varying the DM lifetime, the predicted flux can interpolate between overproduction (smaller $\tau_{\rm DM}$) and underproduction (effectively indistinguishable from background, for larger $\tau_{\rm DM}$), naturally accommodating the KM3NeT detection within a narrow but allowed range of DM lifetime. In the right panel of Fig 4 we show the flux for $M_{\rm DM}=4.5\times 10^{8}$ GeV, and $\tau_{\rm DM}=3.46\times 10^{29}$ s (light blue, solid line) which corresponds to a just $0.131\sigma$ tension with the null detection at IceCube and Auger. We show flux for $\tau_{\rm DM}=1.2\times 10^{30}$ with blue dashed line, which corresponds to a tension of $1.11\sigma$ (This is the same benchmark as shown in the left panel of Fig 1.). This parametric flexibility allows the SHDM scenario to explain the KM3-230213A event without conflicting with IceCube or Auger constraints, thereby providing a competitive and testable new physics interpretation.